OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write a system of two equations with the variables m and f in each to represent this problem.

OpenStudy (anonymous):

Write this out as equations with m=membership fee and f=session fee, so you would have m+5f=25 and m+10f=40 isolate m or f in one of them.... me is clearly easier isolating m in the second equation gives m=40-10f replace that into the first equation to get 40-10f+5f=25 and solve for f and then for m you should get an answer of m=10 and f=3

OpenStudy (anonymous):

are you serious that was easier than my older sister